# Lipschitz-Guided Design of Interpolation Schedules in Generative Models

### v1

@@ 4.3. Invariant distributions of stochastic Allen-Cahn
We consider an infinite-dimensional probability measure defined over continuous functions on the unit interval $[0,1]$, formally proportional to
(4.1) $\exp\left(-\int_{0}^{1}\frac{1}{2}(\partial_{x}u(x))^{2}+V(u(x)){\rm d}x\right),$
where $V(u)=(1-u^{2})^{2}$ is a double-well potential. This is the stationary distribution of the stochastic Allen-Cahn equation
(4.2) $\partial_{t}u=\partial_{xx}u-V^{\prime}(u)+\sqrt{2}\,\eta\,$,
with natural boundary conditions and space-time white noise $\eta$. The distribution is bimodal, with realizations typically exhibiting rough, approximately constant profiles near $u=\pm 1$. We discretize using finite differences on $N$ equidistributed points, yielding an $N$-dimensional distribution. In the interpolant, we sample $x_{1}$ from this distribution using ensemble MCMC algorithms [7]; $z$ is chosen as white noise. We train an ODE generative model and compare energy spectra between true and generated distributions based on different interpolation schedules. The designed schedule is obtained by simply considering the covariance of the Gaussian measure part $\exp(-\int_{0}^{1}\frac{1}{2}(\partial_{x}u(x))^{2}\,{\rm d}x)$ and applying the optimal avg-Lip2 from the Gaussian case (3.11).
@@
Figure 4 demonstrates that for this mildly non-Gaussian behaved distribution, the designed schedule achieves superior accuracy that remains robust across resolutions, unlike linear schedules.
@@ 4.4. Invariant distributions of stochastic Navier-Stokes
Finally, we consider invariant distributions of stochastically forced Navier-Stokes equations on the torus $\mathbb{T}^{2}=[0,2\pi]^{2}$. Using vorticity formulation:
@@
where $v=\nabla^{\perp}\psi=(-\partial_{y}\psi,\partial_{x}\psi)$ is the velocity field from stream function $\psi$ satisfying $\Delta\psi=\omega$. We use parameters $\nu=10^{-3}$, $\alpha=0.1$, $\varepsilon=1$, and white-in-time forcing $\mathrm{d}\eta$ on finite Fourier modes, following [8]. The system is ergodic with a unique invariant measure [18].
We generate data for $x_{1}$ by long-time simulation on a fine grid and use white noise for $z$ in the interpolant. For the designed schedule, we observe that at resolution $128\times 128$, the enstrophy spectrum shows $\sim 10^{-4}$ energy at frequency $k=2^{6}$. We apply the schedule from (3.11) with $\lambda^{*}=10^{-5}$. Figure 5 demonstrates that with 10 RK4 steps, the designed schedule produces superior samples with more accurate spectra. Despite the highly non-Gaussian nature of this distribution, schedules optimized for the Gaussian case can still be used to improve fine-scale accuracy.
@@ 5. Conclusions
In this paper, we studied the design of interpolation schedules in flow and diffusion-based generative models within the stochastic interpolants framework. We revealed a fundamental paradox: while all scalar interpolation schedules achieve identical statistical efficiency under KL divergence in path space after optimal diffusion tuning, their numerical efficiency can differ dramatically. This statistical equivalence result implies that scalar schedule optimization is inherently limited, and future breakthroughs likely require exploring matrix-valued or nonlinear schedules that could break this equivalence barrier.
To exploit the numerical differences among statistically equivalent scalar schedules, we proposed optimizing averaged squared Lipschitzness of the drift field—a criterion that favors schedules requiring fewer integration steps, contrasting with kinetic energy minimization in optimal transport approaches. Our analytical results demonstrate exponential improvements in Lipschitz constants for Gaussian distributions and reduced mode collapse for mixtures. These insights, derived from simple analytical cases, successfully transfer to complex high-dimensional invariant distributions of stochastic Allen-Cahn and Navier-Stokes equations, achieving robust performance across resolutions.
Our scalar schedule optimization demonstrates meaningful practical improvements across diverse applications, from Gaussian distributions to complex stochastic PDEs. However, the statistical equivalence of scalar schedules points toward matrix-valued interpolation schedules as a natural next step that could unlock significantly greater performance gains. In future work, we plan to explore matrix-valued schedules that can adapt individually to different eigenvalue scales, building on the theoretical foundation established here. Additionally, incorporating physics-informed nonlinear schedules and exploring alternative statistical criteria represent promising avenues for further advancing generative model efficiency.
@@ Appendix A Sketch of Derivations for Stochastic Interpolants
$\rho_{t}(x)\propto\int_{\mathbb{R}^{d}}\rho_{t}(y)\exp(-\frac{\|x-y\|_{2}^{2}}{2\alpha_{t}^{2}}){\rm d}y\,.$
@@
$\nabla\log\rho_{t}(x)=\frac{1}{\int_{\mathbb{R}^{d}}\rho_{t}(y)\exp(-\frac{\|x-y\|_{2}^{2}}{2\alpha_{t}^{2}}){\rm d}y}\int_{\mathbb{R}^{d}}(-\frac{x-y}{\alpha_{t}^{2}})\rho_{t}(y)\exp(-\frac{\|x-y\|_{2}^{2}}{2\alpha_{t}^{2}}){\rm d}y\,.$
@@
$\frac{1}{\int_{\mathbb{R}^{d}}\rho_{t}(y)\exp(-\frac{\|x-y\|_{2}^{2}}{2\alpha_{t}^{2}}){\rm d}y}\rho_{t}(y)\exp(-\frac{\|x-y\|_{2}^{2}}{2\alpha_{t}^{2}})$

### v2

@@ 4.3. Invariant distributions of stochastic Allen-Cahn
We consider the invariant distribution of the stochastic Allen–Cahn equation on $[0,1]$, formally given by
$\exp\!\left(-\int_{0}^{1}\frac{1}{2}(\partial_{x}u(x))^{2}+V(u(x))\,\mathrm{d}x\right),\qquad V(u)=(1-u^{2})^{2}.$
This distribution is bimodal and moderately non-Gaussian, with samples concentrating near $u=\pm 1$. We discretize the equation using finite differences on $N$ equispaced grid points, yielding an $N$-dimensional target distribution.
Samples $x_{1}$ from the invariant distribution are generated using ensemble MCMC methods [chen2025new], while $z$ is sampled from spatial white noise. We train an ODE-based generative model and compare the energy spectra of samples generated using different interpolation schedules.
The designed schedule is constructed by considering the covariance of the Gaussian reference measure $\exp(-\int_{0}^{1}\tfrac{1}{2}(\partial_{x}u)^{2}\,\mathrm{d}x)$ and applying the Gaussian avg-Lip2-optimal schedule (3.11). Figure 4 shows that this schedule yields consistently more accurate energy spectra than the linear schedule and that this improvement is robust across spatial resolutions.
@@ 4.4. Invariant distributions of stochastic Navier-Stokes
Finally, we consider invariant distributions of the two-dimensional stochastically forced Navier–Stokes equations on the torus $\mathbb{T}^{2}=[0,2\pi]^{2}$. Using the vorticity formulation, the dynamics are given by
@@
where $v=\nabla^{\perp}\psi=(-\partial_{y}\psi,\partial_{x}\psi)$ is the incompressible velocity field associated with the stream function $\psi$ satisfying $\Delta\psi=\omega$. We fix the parameters to $\nu=10^{-3}$, $\alpha=0.1$, and $\varepsilon=1$. The stochastic forcing $\eta$ is white in time and acts on a finite number of low-frequency Fourier modes, following the setup in [chen2024probabilistic]. Under these choices, the system is ergodic and admits a unique invariant probability measure [hairer2006ergodicity].
Samples from the invariant distribution are generated via long-time numerical simulation of (4.1) using a pseudo-spectral method with standard de-aliasing. After an initial burn-in period sufficient for equilibration, vorticity snapshots are collected at regular time intervals to form an approximately independent dataset of samples. In the stochastic interpolant framework, samples $x_{1}$ are drawn from this empirical invariant distribution, while the initial condition $z$ is sampled from spatial white noise.
We train an ODE-based generative model to transport $z$ to $x_{1}$ over unit time using a learned drift field. The drift is parameterized by a UNet architecture, and training follows the same procedure as in previous sections.
To investigate the effect of interpolation schedules, we compare the standard linear schedule $\beta_{t}=t$ with a Lipschitz-optimized schedule motivated by the Gaussian analysis in Section 3.5. Specifically, we apply the schedule in (3.11) with $\lambda^{\star}=10^{-5}$. This choice is motivated by the empirical observation that, at resolution $128\times 128$, the enstrophy spectrum of the invariant distribution has magnitude on the order of $10^{-4}$ at frequency $k=2^{6}$, while white noise has magnitude around $10$ there, so take $\lambda^{\star}=10^{-4}/10=10^{-5}$.
The ODE is solved using a fixed-step fourth-order Runge–Kutta (RK4) method with a small number of uniform time steps. Figure 5 shows that with only 10 RK4 steps, the designed schedule produces samples with significantly more accurate enstrophy spectra than the linear schedule. Despite the highly non-Gaussian nature of the invariant measure, schedules optimized for the Gaussian case still yield substantial improvements in fine-scale accuracy.
@@ 5. Conclusions
In this paper, we investigated the design of interpolation schedules in flow and diffusion-based generative models within the stochastic interpolants framework. We showed that all scalar interpolation schedules are statistically equivalent under the Kullback–Leibler divergence in path space after optimal tuning of diffusion coefficients, yet can exhibit markedly different numerical behavior. This equivalence result reveals a fundamental limitation of scalar schedule optimization from a statistical perspective and suggests that breaking this barrier will require more expressive choices, such as matrix-valued, nonlinear, or instance-dependent schedules.
To leverage the numerical differences among statistically equivalent scalar schedules, we proposed minimizing the averaged squared Lipschitzness of the drift field as a principled criterion for schedule design. Unlike kinetic energy minimization in optimal transport, this criterion directly targets numerical regularity and favors dynamics that can be accurately integrated with few time steps.
Our analytical results show exponential improvements in Lipschitz constants for Gaussian targets and substantial mitigation of mode collapse for Gaussian mixtures. Importantly, these insights, derived from analytically tractable settings, transfer effectively to complex high-dimensional invariant measures arising from stochastic Allen–Cahn and Navier–Stokes equations, where they yield robust accuracy improvements across spatial resolutions.
Overall, our results demonstrate that carefully designed scalar schedules can deliver meaningful practical gains across a wide range of generative modeling tasks, including challenging stochastic PDE settings. At the same time, the statistical equivalence of scalar schedules highlights the need to move beyond this class in order to achieve further advances. Exploring richer interpolation mechanisms and alternative statistical criteria, as well as understanding their interaction with numerical discretization and physical structure preservation, represents a promising direction for future work in generative modeling for scientific machine learning.
@@ Appendix A Sketch of Derivations for Stochastic Interpolants
$\rho_{t}(x)\propto\int_{\mathbb{R}^{d}}q_{t}(y)\exp(-\frac{\|x-y\|_{2}^{2}}{2\alpha_{t}^{2}}){\rm d}y\,.$
@@
$\nabla\log\rho_{t}(x)=\frac{1}{\int_{\mathbb{R}^{d}}q_{t}(y)\exp(-\frac{\|x-y\|_{2}^{2}}{2\alpha_{t}^{2}}){\rm d}y}\int_{\mathbb{R}^{d}}(-\frac{x-y}{\alpha_{t}^{2}})q_{t}(y)\exp(-\frac{\|x-y\|_{2}^{2}}{2\alpha_{t}^{2}}){\rm d}y\,.$
@@
$\frac{1}{\int_{\mathbb{R}^{d}}q_{t}(y)\exp(-\frac{\|x-y\|_{2}^{2}}{2\alpha_{t}^{2}}){\rm d}y}q_{t}(y)\exp(-\frac{\|x-y\|_{2}^{2}}{2\alpha_{t}^{2}})$